OpenStudy (kittiwitti1):

A 200kg crate is pushed horizontally w/ 700N force. Coefficient of friction is 0.20, what's the crate's acceleration?

OpenStudy (rajat97):

newtons second law

OpenStudy (rajat97):

i am uploading a drawing

OpenStudy (anonymous):

5.462

OpenStudy (kittiwitti1):

the answer is 1.5 m/s/s

OpenStudy (anonymous):

hmm

OpenStudy (kittiwitti1):

this thing is due in 30 minutes ._.

OpenStudy (kittiwitti1):

I have 3 more problems left...

OpenStudy (rajat97):

 total force = f-friction here f=700n and friction=(fric. coeff)xnormal force total force = 700-400=300N now acceleration=total force/mass=300/200=1.5 and you are right